It says solve the proportion
9/4 = (2x+1)/(x+2)

Answer by AnlytcPhil(1810)

It says solve the proportion 
9/4 = (2x+1)/(x+2) 

A proportion is an equation of the form: 
 
"FRACTION = FRACTION"

       9     2x+1
      ——— = ——————
       4      x+2

A proportion is solved by cross multiplying:

LEFT NUMERATOR times RIGHT DENOMINATOR equals 
               LEFT DENOMINATOR times RIGHT NUMERATOR

9 times (x+2) equals 4 times (2x+1)

   9(x+2) = 4(2x+1)

  9x + 18 = 8x + 4

        x = -14

Edwin
AnlytcPhil@aol.com